1. write an equation y=kz
when x=2 y=6
k=
equation=
find y if x=4

Answer by anantha(86)

sol:
let us take k=x
substitute this value in y=kz
y=xz
substitute the values x=2 and y=6 in y=xz
6=2*z
dividing by 2 on both sides
6/2=2*z/2
3=z
substitute these values in y=kz
6=k*3
dividing by 3 on both sides
6/3=k*3/3
k=2
equation y=2z
when x=4
y=xz
y=4*z
6=4z
dividing by 4 on both sides
6/4=4z/4
3/2=z
y=xz
y=4*3/2
y=6